# Large Language Models for Software Engineering: A Systematic Literature Review

### v5

@@ Abstract.
Large Language Models (LLMs) have significantly impacted numerous domains, including Software Engineering (SE). Many recent publications have explored LLMs applied to various SE tasks. Nevertheless, a comprehensive understanding of the application, effects, and possible limitations of LLMs on SE is still in its early stages. To bridge this gap, we conducted a systematic literature review on LLM4SE, with a particular focus on understanding how LLMs can be exploited to optimize processes and outcomes. We collect and analyze 229 research papers from 2017 to 2023 to answer four key research questions (RQs). In RQ1, we categorize different LLMs that have been employed in SE tasks, characterizing their distinctive features and uses. In RQ2, we analyze the methods used in data collection, preprocessing, and application, highlighting the role of well-curated datasets for successful LLM for SE implementation. RQ3 investigates the strategies employed to optimize and evaluate the performance of LLMs in SE. Finally, RQ4 examines the specific SE tasks where LLMs have shown success to date, illustrating their practical contributions to the field. From the answers to these RQs, we discuss the current state-of-the-art and trends, identifying gaps in existing research, and flagging promising areas for future study.
@@
copyright: acmcopyrightjournalyear: 2023doi: XXXXXXX.XXXXXXXccs: General and reference Surveys and overviewsccs: Software and its engineering Software development techniquesccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
In the field of language processing, traditional Language Models (LMs) have been foundational elements, establishing a basis for text generation and understanding (Moore and Lewis, 2010). Increased computational power, advanced machine learning techniques, and access to very large-scale data have led to a significant transition into the emergence of Large Language Models (LLMs) (Zan et al., 2023b; Zhao et al., 2023c). Equipped with expansive and diverse training data, these models have demonstrated an impressive ability to simulate human linguistic capabilities, leading to a sea of changes across multiple domains. With their capacity to learn from massive corpora and generate plausible text, LLMs are blurring the line between human and machine-produced language. They have provided researchers and engineers alike with a powerful tool to explore the complexity and richness of human communication, consequently sparking a transformational period in the field of language processing and beyond.
@@
Simultaneously, researchers have embarked on a series of research activities regarding LLM-related works, where a number of literature reviews or survey papers have been produced (Fan et al., 2023c; Chang et al., 2023; Fan et al., 2023b; Zhao et al., 2023c; Yang et al., 2023a). Table 1 summarises some of these. However, these related studies have limitations. They either focus narrowly on a single SE scope, such as the application of LLMs in software testing (Wang et al., 2023a) and natural-language-to-code (NL2Code) tasks (Zan et al., 2023b), or they are primarily centered on Machine Learning (ML) or Deep Learning (DL) models (Wang et al., 2022a; Yang et al., 2022b), overlooking more advanced and recently emerged LLM applications, such as ChatGPT (OpenAI, 2022a), which are increasingly finding applications within the SE field (Tian et al., 2023b; White et al., 2023b; Lubowitz, 2023; Sridhara et al., 2023). Alternatively, they merely offer a preliminary exploration of the performance of LLMs in various SE tasks through empirical experiments, without conducting a systematic literature survey (Zhao et al., 2023c; Ma et al., 2023a; Sridhara et al., 2023; Xu et al., 2022; Dou et al., 2023; Yuan et al., 2023a). The integration of LLMs within SE is undoubtedly a complex endeavor, requiring key considerations including the choice of the right model, comprehension of the unique features of different LLMs, devising pre-training and fine-tuning strategies, handling of data, evaluation of outcomes, and surmounting implementation challenges (Zan et al., 2023b). Despite the burgeoning interest and ongoing explorations in the field, a detailed and systematic review of LLMs’ application in SE has been notably absent in the current literature. This gap signifies a need for understanding the relationship between LLMs and SE. In response, our research aims to bridge this gap, providing valuable insights to the community.
@@
Zan et al. (Zan et al., 2023b) 2023 LLM (12M+) NL2Code $\times$ 2020-2023 Not specified
Zhao et al. (Zhao et al., 2023c) 2023 LLM (10B+) Beyond SE scope $\times$ - Not specified
Fan et al. (Fan et al., 2023b) 2023 LLM Beyond SE scope $\times$ 2017-2023 5,752
Wang et al. (Wang et al., 2023a) 2023 LLM (117M+) Software testing 2019-2023 52
@@
In this paper, we conduct a systematic literature review on the utilization of LLMs in SE (LLM4SE). By mapping the current state-of-the-art, pinpointing the key strengths, weaknesses, and gaps in the existing LLM4SE literature, and proposing potential avenues for future research, our review aims to provide researchers and practitioners with a thorough guide to the convergence of LLMs and SE. We anticipate that our findings will be instrumental in guiding future inquiries and advancements in this rapidly evolving field. This work makes the following key contributions:
  • We are the first to present a comprehensive systematic literature review based on 229 papers published between 2017 and 2023 that focus on the use of LLM-based solutions to address SE challenges. We conducted a detailed analysis of the selected papers based on publication trends, distribution of publication venues, etc.
@@
  • We discuss optimizers used for LLM4SE tasks, including parameter and learning rate optimization, prevalent prompt engineering techniques, and commonly employed evaluation metrics.
  • We describe the key applications of LLM4SE encompassing a diverse range of 55 specific SE tasks, grouped into six core SE activities – requirements engineering, software design, software development, software quality assurance, software maintenance, and software management.
@@
Section 2 presents our research questions (RQs) and elaborates on our systematic literature review (SLR) methodology. The succeeding Sections 3 to 6 are devoted to answering each of these RQs individually. Section 7 discloses the potential threats to the validity of our study. Section 8 discusses the challenges yet to be overcome when employing LLMs to solve SE tasks and highlights promising opportunities and directions for future research. Section 9 concludes the whole paper.
@@ 2. Approach
This systematic literature review (SLR) follows the methodology proposed by Kitchenham et al. (Kitchenham et al., 2007, 2022), used in most other SE-related SLRs (Li et al., 2017; Wang et al., 2022a; Ramirez et al., 2018; Liu et al., 2022). Following the guidelines provided by Kitchenham et al., our methodology included three main steps: planning the review (i.e., Section 2.1, 2.2), conducting the review (i.e., Section 2.3, 2.4), and analyzing the basic review results (i.e, Section 2.5).
@@ 2.1. Research Questions
To provide a comprehensive overview of the LLM4SE field, it is important to fully comprehend how these models are currently being applied in SE, the challenges they face, and their potential future research directions in SE. Thus, we aim to provide a systematic literature review of the application of LLMs to software engineering. This study thus aims to answer the following research questions:
  • RQ1: What LLMs have been employed to date to solve SE tasks?
  • RQ2: How are SE-related datasets collected, preprocessed, and used in LLMs?
  • RQ3: What techniques are used to optimize and evaluate LLM4SE?
  • RQ4: What SE tasks have been effectively addressed to date using LLM4SE?
@@ 2.2.1. Search Items
During the manual search, we selected six of the top SE conferences and journals (i.e., ICSE, ESEC/FSE, ASE, ISSTA, TOSEM, and TSE, as shown in Table 2) and searched for papers that applied LLM4SE. We systematically crawled a list comprising 4,618 published papers from the top venues. Following automated scanning via scripts, we manually verified and identified 51 papers that were relevant to our research objectives. These 51 relevant papers formed the basis for constructing the Quasi-Gold Standard (QGS). Our search string should combine two sets of keywords: one pertaining to SE tasks, and the other related to LLMs. Only if the paper contains both types of keywords there is a higher probability that it is the paper we need. The complete set of search keywords is as follows:
  • Keywords related to SE tasks: Software Engineering, Software Development, Program*, Software Testing, Software Mainten*, SE, Software Lifecycle, Software Design*, Code representation, Code generation, Code comment generation, Code search, Code localization, Code completion, Code summarization, Method name generation, Bug detection, Bug localization, Vulnerability detection, Testing techniques, Test case generation, Program analysis, Bug classification, Defect prediction, Program repair, Code clone detection, Bug report, Software quality evaluation, SATD detection, Code smell detection, Compiled-related, Code review, Software classification, Code classification, Code change, Incident detection, Requirement extraction, Requirement traceability, Requirement validation, Effort cost prediction, Mining GitHub/Github mining, Mining SO (Stack Overflow)/SO mining, Mining app/App mining, Mining tag/Tag mining, Developer-based mining
@@ 2.2.2. Search Datasets
After determining the search string, we conducted an automated search across seven widely used databases, which are capable of covering all published or latest papers. Given that the first paper about the Transformer architecture (Vaswani et al., 2017), which forms the basis for LLMs, was published in 2017, we focused our search on papers published from that year onward [^1] The cut-off date for the paper collection process of this version is August 1st, 2023.. Two authors independently performed the automated search, and the search results from each database were merged and deduplicated. Specifically, we obtained 1,089 papers from IEEE Xplore, 5,433 papers from the ACM Digital Library, 44,158 papers from ScienceDirect, 31,894 papers from Web of Science, 70,946 papers from Springer, 8,254 papers from arXiv, and 2,592 papers from DBLP.
@@ 2.3.1. Study Inclusion and Exclusion Criteria
1) The paper claims that an LLM is used
2) The paper claims that the study involves an SE task
3) The paper with accessible full text
@@
1) Short papers whose number of pages is less than 8
@@
2) Duplicate papers or similar studies with different versions from the same authors
@@
3) Studies belonging to books, thesis, monographs, keynotes, panels, or venues not
executing a full peer-review process
4) Tool demos and editorials
5) The paper is published in a workshop or a doctoral symposium
6) The paper is a grey publication, e.g., a technical report or thesis
7) Non-English written literature
8) Literature mentioning the use of LLMs without describing the employed techniques
Based on our search strategy, we initially obtained 164,366 papers that potentially relate to our research. Next, we needed to further evaluate the relevance of these papers based on inclusion and exclusion criteria, as shown in Table 3, so that the selected papers can directly address our research questions. The paper selection process, as illustrated in Fig. 1, consists of six phases.
@@
In the first phase, we conducted automated filtering to exclude papers with less than 8 pages (Bashroush et al., 2017; Wang et al., 2022a) (Exclusion criteria 1), reducing the number of papers to 63,404. In the second phase, we examined the titles, abstracts, and keywords of the papers to identify those that include relevant LLM-related keywords. We then expanded the search scope to avoid missing relevant papers, including ML, DL, and other related keywords that may not directly correspond to LLM. The purpose of this phase is to narrow down the scope and filter out papers directly related to LLM (Inclusion criteria 1). Papers that are filtered out in this phase are then manually reviewed in the fifth phase. Additionally, we excluded 235 non-English written literature (Exclusion criteria 7). After the second phase, the number of papers was reduced to 4,341.
@@
The third phase involves identifying the venues of the papers (Exclusion criteria 3). We extracted publication information such as “journal”, “URL”, “DOI”, and “series” to determine the publication sources. For papers from arXiv in 2022 and 2023, we chose to retain them, considering that this field is emerging and many works are in the process of submission. Although these papers did not undergo peer review, we have a quality assessment process to eliminate papers with low quality, ensuring the overall quality of this systematic literature review (SLR). This step resulted in 632 papers.
@@
In the fourth phase, we merged and deduplicated the remaining papers from the seven databases and the manually searched paper list (Exclusion criteria 2), resulting in 548 papers. We then reviewed the full texts of the papers and excluded 176 papers that were grey publications or were published in workshops or doctoral symposiums (Exclusion criteria 4, 5, 6). By further assessing the quality of the papers, we identified 218 papers directly relevant to our research. This phase primarily involved excluding papers that mentioned LLMs but did not directly apply them, such as papers that only discussed LLMs in future work or focused on evaluating the performance of LLM-enabled tools (Wang et al., 2023a) (Exclusion criteria 8). For SLR, survey, and review papers, we have retained them and will assess their content during the quality assessment phase to determine their relevance to our research.

### v6

@@ Abstract.
Large Language Models (LLMs) have significantly impacted numerous domains, including Software Engineering (SE). Many recent publications have explored LLMs applied to various SE tasks. Nevertheless, a comprehensive understanding of the application, effects, and possible limitations of LLMs on SE is still in its early stages. To bridge this gap, we conducted a systematic literature review (SLR) on LLM4SE, with a particular focus on understanding how LLMs can be exploited to optimize processes and outcomes. We select and analyze 395 research papers from January 2017 to January 2024 to answer four key research questions (RQs). In RQ1, we categorize different LLMs that have been employed in SE tasks, characterizing their distinctive features and uses. In RQ2, we analyze the methods used in data collection, preprocessing, and application, highlighting the role of well-curated datasets for successful LLM for SE implementation. RQ3 investigates the strategies employed to optimize and evaluate the performance of LLMs in SE. Finally, RQ4 examines the specific SE tasks where LLMs have shown success to date, illustrating their practical contributions to the field. From the answers to these RQs, we discuss the current state-of-the-art and trends, identifying gaps in existing research, and flagging promising areas for future study. Our artifacts are publicly available at https://github.com/xinyi-hou/LLM4SE_SLR.
@@
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: TOSEMjournalvolume: Xjournalnumber: Yarticle: 1publicationmonth: 12ccs: General and reference Surveys and overviewsccs: Software and its engineering Software development techniquesccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
In the field of language processing, traditional Language Models (LMs) have been foundational elements, establishing a basis for text generation and understanding (Moore and Lewis, 2010). Increased computational power, advanced machine learning techniques, and access to very large-scale data have led to a significant transition into the emergence of Large Language Models (LLMs) (Zan et al., 2023b; Zhao et al., 2023d). Equipped with expansive and diverse training data, these models have demonstrated an impressive ability to simulate human linguistic capabilities, leading to a sea of changes across multiple domains. With their capacity to learn from massive corpora and generate plausible text, LLMs are blurring the line between human and machine-produced language. They have provided researchers and engineers alike with a powerful tool to explore the complexity and richness of human communication, consequently sparking a transformational period in the field of language processing and beyond.
@@
Simultaneously, researchers have embarked on a series of research activities regarding LLM-related works, where a number of literature reviews or survey papers have been produced (Fan et al., 2023c; Chang et al., 2023; Yang et al., 2023b). Table 1 summarises some of these. However, these related studies have limitations. They either focus narrowly on a single SE scope, such as the application of LLMs in software testing (Wang et al., 2023c) and natural-language-to-code (NL2Code) tasks (Zan et al., 2023b), or they are primarily centered on Machine Learning (ML) or Deep Learning (DL) models (Wang et al., 2022a; Watson et al., 2022; Yang et al., 2022b), overlooking more advanced and recently emerged LLM applications, such as ChatGPT (OpenAI, 2022a), which are increasingly finding applications within the SE field (Tian et al., 2023b; White et al., 2023b; Lubowitz, 2023; Sridhara et al., 2023). Alternatively, they merely offer a preliminary exploration of the performance of LLMs in various SE tasks through empirical experiments (Ma et al., 2023a; Sridhara et al., 2023; Xu et al., 2022; Dou et al., 2023; Yuan et al., 2023a), or analyze existing partially relevant studies to reveal the challenges in this field (Fan et al., 2023b) without conducting a systematic literature survey. Furthermore, some works have investigated the application of Code LLMs in SE (Zhang et al., 2023b; Zheng et al., 2023a), yet have not fully considered general LLMs like ChatGPT and LLaMA (Touvron et al., 2023a), which are also widely applied to various SE tasks (Huang et al., 2023a; Shapkin et al., 2023; Pan et al., 2023a; Yan et al., 2023a). The integration of LLMs within SE is undoubtedly a complex endeavor, requiring key considerations including the choice of the right model, comprehension of the unique features of different LLMs, devising pre-training and fine-tuning strategies, handling of data, evaluation of outcomes, and surmounting implementation challenges (Zan et al., 2023b). Despite the burgeoning interest and ongoing explorations in the field, a detailed and systematic review of LLMs’ application in SE has been notably absent in the current literature. This gap signifies a need for understanding the relationship between LLMs and SE. In response, our research aims to bridge this gap, providing valuable insights to the community.
@@
Zhang et al. (Zhang et al., 2023b) 2023 Code LLM Automated program repair 2017-2023 185
Zheng et al. (Zheng et al., 2023a) 2023 Code LLM General SE scope 2021-2023 149
Fan et al. (Fan et al., 2023b) 2023 LLM General SE scope $\times$× - Not specified
Zan et al. (Zan et al., 2023b) 2023 LLM (12M+) NL2Code $\times$× 2020-2023 Not specified
Wang et al. (Wang et al., 2023c) 2023 LLM (117M+) Software testing 2019-2023 52
@@
Watson et al. (Watson et al., 2022) 2022 DL General SE scope 2009-2019 128
Our work 2024 LLM General SE scope 2017-2024 395
@@
In this paper, we conduct an SLR on the utilization of LLMs in SE (LLM4SE). By mapping the current state-of-the-art, pinpointing the key strengths, weaknesses, and gaps in the existing LLM4SE literature, and proposing potential avenues for future research, our review aims to provide researchers and practitioners with a thorough guide to the convergence of LLMs and SE. We anticipate that our findings will be instrumental in guiding future inquiries and advancements in this rapidly evolving field. This work makes the following key contributions:
  • We are the first to present a comprehensive SLR on 395 papers published between January 2017 and January 2024 that focus on the use of LLM-based solutions to address SE challenges. We conducted a detailed analysis of the selected papers based on publication trends, distribution of publication venues, etc.
@@
  • We discuss optimizers used for LLM4SE tasks, including tuning techniques, prevalent prompt engineering techniques, and commonly employed evaluation metrics.
  • We describe the key applications of LLM4SE encompassing a diverse range of 85 specific SE tasks, grouped into six core SE activities – requirements engineering, software design, software development, software quality assurance, software maintenance, and software management.
@@
Section 2 presents our research questions (RQs) and elaborates on our SLR methodology. The succeeding Sections 3 to 6 are devoted to answering each of these RQs individually. Section 7 discloses the potential threats to the validity of our study. Section 8 discusses the challenges yet to be overcome when employing LLMs to solve SE tasks and highlights promising opportunities and directions for future research. Section 9 concludes the whole paper.
@@ 2. Approach
This SLR follows the methodology proposed by Kitchenham et al. (Kitchenham et al., 2007, 2022), used in most other SE-related SLRs (Li et al., 2017; Wang et al., 2022a; Ramirez et al., 2018; Liu et al., 2022b). Following the guidelines provided by Kitchenham et al., our methodology included three main steps: planning the review (i.e., Section 2.1, 2.2), conducting the review (i.e., Section 2.3, 2.4), and analyzing the basic review results (i.e, Section 2.5).
@@ 2.1. Research Questions
To provide a comprehensive overview of the LLM4SE field, it is important to fully comprehend how these models are currently being applied in SE, the challenges they face, and their potential future research directions in SE. Thus, we aim to provide an SLR of the application of LLMs to software engineering. This study thus aims to answer the following research questions:
RQ1: What LLMs have been employed to date to solve SE tasks? RQ1 is designed to map out the landscape of LLMs applied in the field of SE. It seeks to identify and categorize the various LLM architectures—such as decoder-only, encoder-decoder, and encoder-only models—that have been leveraged to address diverse SE challenges. This RQ aims to provide a comprehensive overview of how these models are being utilized and the implications of their usage in this field.
RQ2: How are SE-related datasets collected, preprocessed, and used in LLMs? RQ2 delves into the methodologies behind the assembly, refinement, and application of datasets in the realm of LLMs for SE tasks. It aims to uncover the strategies for dataset collection, the criteria for dataset selection, and the preprocessing steps essential for making the data conducive for LLM training and application. Additionally, this question seeks to explore the types of data that are most prevalent in SE-related LLM research and how these data types influence the modeling and outcomes.
RQ3: What techniques are used to optimize and evaluate LLM4SE? RQ3 aims to explore the use of different optimization and evaluation techniques specific to LLMs in the context of SE. This includes an investigation into Parameter Efficient Fine-Tuning (PEFT) methods and various prompting techniques that are tailored to enhance LLM performance on SE tasks. Furthermore, this RQ aims to assess the range of evaluation metrics and methodologies employed to gauge the effectiveness and impact of LLMs in SE, providing insights into how these models are fine-tuned and assessed for their utility and efficiency.
RQ4: What SE tasks have been effectively addressed to date using LLM4SE? This RQ aims to identify the SE tasks that have been successfully tackled using LLMs, offering a detailed view of the application spectrum of LLMs in SE. It seeks to identify the specific tasks within SE, such as code generation and program repair, where LLMs have shown significant utility, and to explore the nature and scope of these applications.
@@ 2.2.1. Search Items
During the manual search, we selected six of the top SE conferences and journals (i.e., ICSE, ESEC/FSE, ASE, ISSTA, TOSEM, and TSE, as shown in Table 2) and searched for papers that applied LLM4SE. We systematically crawled a list comprising 4,618 published papers from the top venues. Following automated scanning via scripts, we manually verified and identified 51 papers that were relevant to our research objectives. These 51 relevant papers formed the basis for constructing the Quasi-Gold Standard (QGS). Our search string should combine two sets of keywords: one pertaining to SE tasks, and the other related to LLMs. Only if the paper contains both types of keywords, there is a higher probability that it is the paper we need. The complete set of search keywords is as follows:
  • Keywords related to SE tasks: Software Engineering, Software Development, Program* [^1] The * symbol serves as a wildcard, representing any characters or character sequence. For example, “Program*” can match “Program”, “Programming”, “Programmer”, and so on., Software Testing, Software Mainten*, SE, Software Lifecycle, Software Design*, Code representation, Code generation, Code comment generation, Code search, Code localization, Code completion, Code summarization, Method name generation, Bug detection, Bug localization, Vulnerability detection, Testing techniques, Test case generation, Program analysis, Bug classification, Defect prediction, Program repair, Code clone detection, Bug report, Software quality evaluation, SATD detection, Code smell detection, Compiled-related, Code review, Software classification, Code classification, Code change, Incident detection, Requirement extraction, Requirement traceability, Requirement validation, Effort cost prediction, Mining GitHub/Github mining, Mining SO (Stack Overflow)/SO mining, Mining app/App mining, Mining tag/Tag mining, Developer-based mining
@@ 2.2.2. Search Datasets
After determining the search string, we conducted an automated search across seven widely used databases, which are capable of covering all published or latest papers. Given that the first paper about the Transformer architecture (Vaswani et al., 2017), which forms the basis for LLMs, was published in 2017, we focused our search on papers published from that year onward [^2] The cut-off date for the paper collection process of this version is January 31, 2024.. Two authors independently performed the automated search, and the search results from each database were merged and deduplicated. Specifically, we obtained 1,192 papers from IEEE Xplore, 10,445 papers from the ACM Digital Library, 62,290 papers from ScienceDirect, 42,166 papers from Web of Science, 85,671 papers from Springer, 9,966 papers from arXiv, and 4,035 papers from DBLP.
@@ 2.3.1. Study Inclusion and Exclusion Criteria
1) The paper claims that an LLM is used.
2) The paper claims that the study involves an SE task.
3) The paper with accessible full text.
@@
1) Short papers whose number of pages is less than 8.
@@
2) Duplicate papers or similar studies with different versions from the same authors.
@@
3) Studies belonging to books, thesis, monographs, keynotes, panels, or venues not executing a full
peer-review process.
4) Tool demos and editorials.
5) The paper is published in a workshop or a doctoral symposium.
6) The paper is a grey publication, e.g., a technical report or thesis.
7) Non-English written literature.
8) The paper mentions the use of LLMs without describing the employed techniques.
9) The paper leverages SE methods to enhance LLMs, rather than focusing on using LLMs for SE tasks.
@@
Based on our search strategy, we initially obtained 218,765 papers that potentially relate to our research. Next, we needed to further evaluate the relevance of these papers based on inclusion and exclusion criteria (To ensure that our inclusion and exclusion criteria were sufficiently objective and rational, we designed these criteria following several state-of-the-art SLR papers (Naveed et al., 2024; Wang et al., 2022a; Watson et al., 2022; Yang et al., 2022b).), as shown in Table 3, so that the selected papers can directly address our research questions. The paper selection process, as illustrated in Fig. 1, consists of six phases. In the first phase, we conducted automated filtering to exclude papers with less than 8 pages (Bashroush et al., 2017; Wang et al., 2022a) (Exclusion criteria 1), reducing the number of papers to 80,611. In the second phase, we examined the titles, abstracts, and keywords of the papers to identify those that include relevant LLM-related keywords. We then expanded the search scope to avoid missing relevant papers, including ML, DL, and other related keywords that may not directly correspond to LLM. The purpose of this phase is to narrow down the scope and filter out papers directly related to LLM (Inclusion criteria 1). Papers that are filtered out in this phase are then manually reviewed in the fifth phase. Additionally, we excluded 448 non-English written literature (Exclusion criteria 7). After the second phase, the number of papers was reduced to 5,078.
@@
The third phase involves identifying the venues of the papers (Exclusion criteria 3). We extracted publication information such as “journal”, “URL”, “DOI”, and “series” to determine the publication sources. For papers from arXiv in 2023 and 2024, we chose to retain them, considering that this field is emerging and many works are in the process of submission. Although these papers did not undergo peer review, we have a quality assessment process to eliminate papers with low quality. This step resulted in 1,172 papers.
@@
In the fourth phase, we merged and deduplicated the remaining papers from the seven databases and the manually searched paper list (Exclusion criteria 2), resulting in 810 papers. We then reviewed the full texts of the papers and excluded 190 papers that were grey publications or were published in workshops or doctoral symposiums (Exclusion criteria 4, 5, 6). By further assessing the quality of the papers, we identified 382 papers directly relevant to our research. This phase primarily involved excluding papers that mentioned LLMs but did not directly apply them, such as papers that only discussed LLMs in future work or focused on evaluating the performance of LLM-enabled tools (Wang et al., 2023c) (Exclusion criteria 8). For systematic views, survey, and review papers, we have retained them and will assess their content during the quality assessment phase to determine their relevance to our research.